OpenStudy (anonymous):

solve using the substitution method 2x+5y=14 -5x+y=19

OpenStudy (anonymous):

Pick one equation. Solve for one variable, then used that formula to substitute for that variable in the other equation. Now you have one equation with one variable, solve for it and you have one solution. Substitute that value in either equation and you're done.

OpenStudy (anonymous):

From second equation: y=19+5x Substitute in first equation: 2x+5(19+5x)=14 2x+95+25x=14 27x=14-95 27x=-81 x=-3 Finally sub in the second one (or the first would work too) -5(-3)+y=19 15+y=19 y=4